OpenStudy (anonymous):

I have a fraction that needs to be in simplest form. This is what It looks like 9 6/2. Please SLOWLY explain how I do this.

OpenStudy (skullpatrol):

$$\huge 9 \frac{6}{2}$$

OpenStudy (skullpatrol):

@Prettyfi13 is this what you have?

OpenStudy (skullpatrol):

This is called a "mixed number" because it is the sum of an integer and a fraction.

OpenStudy (skullpatrol):

$$\huge 9\frac{6}{2}=9+\frac{6}{2}$$ $$\huge \frac{6}{2}=6\div2=3$$ $$\huge \therefore 9\frac{6}{2}=9+3=12$$